# Two tales for a geometric Jensen–Shannon divergence

### v1

@@ 5 Summary and concluding remarks
$\mathrm{JS}_{D}^{M_{\alpha},\beta}(p_{1},p_{2}){:=}\beta\,D\left(p_{1},{(p_{1}p_{2})}_{\alpha}^{M}\right)+(1-\beta)\,D\left(p_{2},{(p_{1}p_{2})}_{\alpha}^{M}\right),\quad(\alpha,\beta)\in(0,1)^{2},$
@@
Then we defined two types of geometric Jensen–Shannon divergence between probability densities: The first type $D_{\mathrm{JS}}^{M}$ requires to normalize $M$-mixtures and relies on the Kullback–Leibler divergence: $D_{\mathrm{JS}}^{M}=\mathrm{JS}_{\mathrm{KL}}^{M_{\frac{1}{2}},\frac{1}{2}}$. The second type $D_{\mathrm{JS}^{+}}^{\tilde{M}}$ does not normalize $M$-mixtures and uses the extended Kullback–Leibler divergence: $D_{\mathrm{JS}^{+}}^{\tilde{M}}=\mathrm{JS}_{{\mathrm{KL}}^{+}}^{M_{\frac{1}{2}},\frac{1}{2}}$. When $M$ is the arithmetic mean $A$, both $M$-JSD types coincide with the ordinary Jensen–Shannon divergence of Eq. 2.
We have shown that that both geometric JSD types can be interpreted as regularized Jensen–Shannon divergences $D_{\mathrm{JS}}$ by additive terms:
$D_{\mathrm{JS}}^{M}(p_{1},p_{2})$ $=$ ${\bf D_{\mathrm{JS}}(p_{1},p_{2})}+D_{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M}),$
$D_{\mathrm{JS}^{+}}^{\tilde{M}}(p_{1},p_{2})$ $=$ $D_{\mathrm{JS}}^{M}(p_{1},p_{2})+Z_{M}(p_{1},p_{2})-\log Z_{M}(p_{1},p_{2})-1,$
$=$ ${\bf D_{\mathrm{JS}}(p_{1},p_{2})}+D_{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M})+Z_{M}(p_{1},p_{2})-\log Z_{M}(p_{1},p_{2})-1,$
where $Z_{M}(p_{1},p_{2})=\int M(p_{1},p_{2})\,{\mathrm{d}\mu}$ is the $M$-mixture normalizer.
We have expressed $D_{\mathrm{JS}}^{G}$ using both Jeffreys $D_{J}$ and Bhattacharrya $D_{B}$ divergences as follows:
$D_{\mathrm{JS}}^{G}(p_{1},p_{2})=\frac{1}{4}\,D_{J}(p_{1},p_{2})-D_{B}(p_{1},p_{2}).$
We obtained both lower and upper bounds on $D_{\mathrm{JS}}^{G}$:
$0\leq D_{\mathrm{JS}}(p_{1},p_{2})\leq D_{\mathrm{JS}}^{G}(p_{1},p_{2})\leq\frac{1}{4}\,D_{J}(p_{1},p_{2}).$
We studied power JSDs by considering the power means and study in the $\pm\infty$ limits, the extended max-JSD and min-JSD: We proved that the extended max-JSD is upper bounded by the total variation distance $D_{\mathrm{TV}}(p_{1},p_{2})=\frac{1}{2}\,\int|p_{1}-p_{2}|\,{\mathrm{d}\mu}$:
$0\leq D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\leq D_{\mathrm{TV}}(p_{1},p_{2}),$
@@
$D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})\geq\frac{1}{4}D_{J}(p_{1},p_{2})-D_{\mathrm{TV}}(p_{1},p_{2}),$
where $D_{J}$ denotes Jeffreys’s divergence: $D_{J}(p_{1},p_{2})=D_{\mathrm{KL}}(p_{1},p_{2})+D_{\mathrm{KL}}(p_{2},p_{1})$.
The advantage of using the extended G-JSD is that we do not need to normalize geometric mixtures while this novel divergence retains the property that it amounts to a regularization of the ordinary Jensen–Shannon divergence by an extra additive gap term.
Finally, we expressed $D_{\mathrm{JS}}^{G}$ (Eq. 40) and $D_{\mathrm{JS}^{+}}^{\tilde{G}}$ (Eq. 41) for exponential families, characterized the gap between these two types of divergences as a function of the cumulant and partition functions, and reported corresponding explicit formula for the multivariate Gaussian (exponential) family. The G-JSD between Gaussian distributions has already been used successfully in many applications [8, 20, 18, 23, 33, 39, 34, 37, 13].
As a final word, let us also mention a recent work that considered the geometric Jensen–Shannon divergence between Gaussian measures in infinite-dimensional Hilbert spaces [32] (with applications to Gaussian processes left for future work).

### v3

@@ 6 Summary and concluding remarks
$D^{\mathrm{JS}}_{M_{\alpha},\beta}(p_{1},p_{2}){:=}\beta\,D\left(p_{1},{(p_{1}p_{2})}_{M_{\alpha}}\right)+(1-\beta)\,D\left(p_{2},{(p_{1}p_{2})}_{M_{\alpha}}\right),\quad(\alpha,\beta)\in(0,1)^{2},$
@@
Then we defined two types of geometric Jensen–Shannon divergence between probability densities: The first type $\mathrm{JS}_{M}$ requires to normalize $M$-mixtures and relies on the Kullback–Leibler divergence: $\mathrm{JS}_{M}={\mathrm{KL}}^{\mathrm{JS}}_{M_{\frac{1}{2}},\frac{1}{2}}$. The second type $\mathrm{JS}^{+}_{\tilde{M}}$ does not normalize $M$-mixtures and uses the extended Kullback–Leibler divergence $\mathrm{KL}$$^{+}$ to take into account unnormalized mixtures: $\mathrm{JS}^{+}_{\tilde{M}}={\mathrm{KL}}^{\mathrm{JS}^{+}}_{{\tilde{M}}_{\frac{1}{2}},\frac{1}{2}}$. When $M$ is the arithmetic mean $A$, both $M$-JSD types coincide with the ordinary Jensen–Shannon divergence of Eq. 2.
We have shown that both $M$-JSD types can be interpreted as regularized Jensen–Shannon divergences $\mathrm{JS}$ with additive terms. Namely, we have:
$\mathrm{JS}_{M}(p_{1},p_{2})$ $=$ ${\mathrm{JS}(p_{1},p_{2})}+{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M}),$
$\mathrm{JS}^{+}_{\tilde{M}}(p_{1},p_{2})$ $=$ $\mathrm{JS}_{M}(p_{1},p_{2})+Z_{M}(p_{1},p_{2})-\log Z_{M}(p_{1},p_{2})-1,$
$=$ ${\mathrm{JS}(p_{1},p_{2})}+{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M})+Z_{M}(p_{1},p_{2})-\log Z_{M}(p_{1},p_{2})-1,$
where $Z_{M}(p_{1},p_{2})=\int M(p_{1},p_{2})\,{\mathrm{d}\mu}$ is the $M$-mixture normalizer. The gap between these two types of M-JSD is
$\Delta_{M}(p_{1},p_{2})$ $=$ $\mathrm{JS}_{{\tilde{M}}}^{+}(p_{1},p_{2})-\mathrm{JS}_{M}(p_{1},p_{2}),$
$=$ $Z_{M}(p_{1},p_{2})-\log Z_{M}(p_{1},p_{2})-1.$
When taking the geometric mean $M=G$, we showed that both G-JSD types can be expressed using the Jeffreys divergence and the Bhattacharyya divergence (or Bhattacharyya coefficient):
$\mathrm{JS}_{G}(p_{1},p_{2})$ $=$ $\frac{1}{4}\,J(p_{1},p_{2})-B(p_{1},p_{2}),$
$\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})$ $=$ $\frac{1}{4}\,J(p_{1},p_{2})+\exp(-B(p_{1},p_{2}))-1,$
$=$ $\frac{1}{4}\,J(p_{1},p_{2})+\mathrm{BC}(p_{1},p_{2})-1.$
Thus the gap between these two types of G-JSD is
$\Delta_{G}(p_{1},p_{2})$ ${:=}$ $\mathrm{JS}_{{\tilde{G}}}^{+}(p_{1},p_{2})-\mathrm{JS}_{G}(p_{1},p_{2}),$
$=$ $\mathrm{BC}(p_{1},p_{2})+B(p_{1},p_{2})-1,$
$=$ $Z_{G}(p_{1},p_{2})-\log Z_{G}(p_{1},p_{2})-1,$
since $Z_{G}(p_{1},p_{2})=\int\sqrt{p_{1}\,p_{2}}{\mathrm{d}\mu}=\mathrm{BC}(p_{1},p_{2})$.
Although the square root of the Jensen–Shannon divergence yields a metric distance, this is not anymore the case for the geometric-JSD and the extended geometric-JSD: We reported counterexamples in Remark 9. Moreover, we have shown that the KL symmetrization $\sqrt{{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{G})}$ is not a metric distance (Remark 13).
We discussed the merit of the extended G-JSD which does not require to normalize the geometric mixture in §5, and showed how to approximate the G-JSD using the projective $\gamma$-divergences [19] for $\gamma=\epsilon$, a small enough value (i.e., $\gamma=\epsilon=10^{-3}$). From the viewpoint of information geometry, the extended G-JSD has been shown to be a $f$-divergence [3] (separable divergence) while the G-JSD is not separable in general because of the normalization of mixtures (with exception of the ordinary JSD which is a $f$-divergence because the arithmetic mixtures do not require normalization).
We studied power JSDs by considering the power means and study in the $\pm\infty$ limits, the extended max-JSD and min-JSD: We proved that the extended max-JSD is upper bounded by the total variation distance $\mathrm{TV}(p_{1},p_{2})=\frac{1}{2}\,\int|p_{1}-p_{2}|\,{\mathrm{d}\mu}$:
$0\leq\mathrm{JS}^{+}_{\widetilde{\max}}(p_{1},p_{2})\leq\mathrm{TV}(p_{1},p_{2}),$
@@
$\mathrm{JS}^{+}_{\widetilde{\min}}(p_{1},p_{2})\geq\frac{1}{4}\,J(p_{1},p_{2})-\mathrm{TV}(p_{1},p_{2}),$
where $J$ denotes Jeffreys’s divergence: $J(p_{1},p_{2})={\mathrm{KL}}(p_{1},p_{2})+{\mathrm{KL}}(p_{2},p_{1})$.
The advantage of using the extended G-JSD is that we do not need to normalize geometric mixtures while this novel divergence is proven to be a $f$-divergence [3] and retains the property that it amounts to a regularization of the ordinary Jensen–Shannon divergence by an extra additive gap term.
Finally, we expressed $\mathrm{JS}_{G}$ (Eq. 42) and $\mathrm{JS}^{+}_{\tilde{G}}$ (Eq. 43) for exponential families, characterized the gap between these two types of divergences as a function of the cumulant and partition functions, and reported corresponding explicit formula for the multivariate Gaussian (exponential) family. The G-JSD between Gaussian distributions has already been used successfully in many applications [16, 31, 35, 48, 56, 50, 54, 23].
Appendix A Notations
Means:
$M_{\alpha}(a,b)$ weighted scalar mean
$M_{\alpha}^{\phi}(a,b)$ weighted quasi-arithmetic scalar mean for generator $\phi(u)$
$A(a,b)$ arithmetic mean
$A_{\alpha}(a,b)$ weighted arithmetic mean
$G_{\alpha}(a,b)$ weighted geometric mean
$G(a,b)$ geometric mean
$P_{\gamma}(a,b)$ power mean with $P_{0}=G$ and $P_{1}=A$
$P_{\gamma,\alpha}(a,b)$ weighted power mean
Densities on measure space $(\mathcal{X},\mathcal{E},\mu)$:
$p,p_{1},p_{2},\ldots$ normalized density
$q,q_{1},q_{2},\ldots$ unnormalized density
$Z(q)$ density normalizer $p=\frac{q}{Z(q)}$
$Z_{M}(p_{1},p_{2})$ normalizer of $M$-mixture ($\alpha=\frac{1}{2}$)
$\hat{Z}_{M}(p_{1},p_{2})$ Monte Carlo estimator of $Z_{M}(p_{1},p_{2})$
$Z_{M,\alpha}(p_{1},p_{2})$ normalizer of weighted $M$-mixture
$(p_{1}p_{2})_{M}$ $M$-mixture
$(p_{1}p_{2})_{M,\alpha}$ weighted $M$-mixture
Dissimilarities, divergences, and distances:
${\mathrm{KL}}(p_{1},p_{2})$ Kullback–Leibler divergence (KLD)
${\mathrm{KL}}^{+}(q_{1},q_{2})$ extended Kullback–Leibler divergence
${\mathrm{KL}}^{*}(p_{1},p_{2})$ reverse Kullback–Leibler divergence
$H^{\times}(p_{1},p_{2})$ cross-entropy
$H(p)$ Shannon discrete or differential entropy
$J(p_{1},p_{2})$ Jeffreys divergence
$\mathrm{TV}(p_{1},p_{2})$ total variation distance
$B(p_{1},p_{2})$ Bhattacharyya “distance” (not metric)
$B_{\alpha}(p_{1},p_{2})$ $\alpha$-skewed Bhattacharyya “distance”
$C(p_{1},p_{2})$ Chernoff information or Chernoff distance
$T(p_{1},p_{2})$ Taneja $T$-divergence
$I_{f}(p_{1},p_{2})$ Ali-Silvey-Csiszár $f$-divergence
$D(p_{1},p_{2})$ arbitrary dissimilarity measure
$D^{*}(p_{1},p_{2})$ reverse dissimilarity measure
$D^{+}(q_{1},q_{2})$ extended dissimilarity measure
${\tilde{D}}(q_{1},q_{2})$ projective dissimilarity measure
${\tilde{D}}_{\gamma}(q_{1},q_{2})$ $\gamma$-divergence
$\hat{D}^{+}(q_{1},q_{2})$ Monte Carlo estimation of dissimilarity $D^{+}$
Jensen–Shannon divergences and generalizations:
$\mathrm{JS}(p_{1},p_{2})$ Jensen–Shannon divergence (JSD)
$\mathrm{JS}_{\alpha,\beta}(p_{1},p_{2})$ $\beta$-weighted $\alpha$-skewed mixture JSD
$\mathrm{JS}_{M}(p_{1},p_{2})$ $M$-JSD for $M$-mixtures
$\mathrm{JS}_{G}(p_{1},p_{2})$ geometric JSD
$\mathrm{JS}_{{\tilde{G}}}(p_{1},p_{2})$ extended geometric JSD
$\mathrm{JS}_{G}^{*}(p_{1},p_{2})$ left-sided geometric JSD (right-sided for $\mathrm{KL}$$^{*}$)
${\mathrm{JS}^{+}}_{\widetilde{\min}}(p_{1},p_{2})$ min-JSD
${\mathrm{JS}^{+}}_{\widetilde{\max}}(p_{1},p_{2})$ max-JSD
$\Delta_{M}(p_{1},p_{2})$ gap between extended and normalized M-JSDs